OpenStudy (anonymous):

help! 2-4+6x7-10+2x5=

OpenStudy (lgbasallote):

multiply everything that can be multiplied first

OpenStudy (anonymous):

soooo pemdas?

OpenStudy (anonymous):

Yup...

OpenStudy (mayankdevnani):

2-4+6*7-10+2*5 =-2+42-10+10 =40

OpenStudy (mayankdevnani):

@Kaitlyn_The_Math_wiz

OpenStudy (anonymous):

@mayankdevnani the answer will be= 2-4+6*7-10+2*5=-2+42-10+10=40

OpenStudy (mayankdevnani):

@pourushgupta so, what my answer????